15-11 Jansken Co. has an ROE = .11 and a beta of 1.20. It plans to maintain indefinitely its traditional payout ratio of 0.35. This year’s earnings were $2.50 per share. This year’s dividend was just paid. The consensus estimate of the coming year’s market return is 10%, and T-bills currently offer 5% return. Find the price at which Jansken should sell using the DDM.
Answer #1

Growth rate = ROE * Retention ratio

Retention Ratio = 1 - Div Payout Ratio

= 1 - 0.35

= 0.65

Growth rate = ROE * Retention ratio

= 0.11 * 0.65

= 7.15%

D1 = EPS* Payout ratio ( 1 + g)

= 2.50 * 35% ( 1 + 0.0715 )

= 0.875 (1.0715)

= $ 0.9376

Ke = Rf + beta ( Rm - Rf)

= 5% + 1.20 ( 10% -5 % )

= 5% + 1.2 (5% )

= 5% +6%

= 11%

Price = D1 / [ Ke - g ]

=0.9376 / [ 11% - 7.15% ]

= 0.9376 / 3.85%

= $ 24.35

Price of Stock is $24.35